OpenStudy (anonymous):

Find the exact value of the following expression. If undefined, state, undefined. sec((cot^(-1)(5/12))

OpenStudy (anonymous):

use the identity \[\sec^2 x=1+\tan^2x\]

OpenStudy (anonymous):

thanks littel confused I keep havign a sin leftover

OpenStudy (anonymous):

pls explain

OpenStudy (anonymous):

there is a right angled triangle with sides 13,5,12 therefore the angle whose cot is 5/12 has a sec of 13/5 Draw the triangle and you will see.

OpenStudy (anonymous):

using the formula if cot = 5/12 then tan = 12/5 so sec^2x = 1 + 12^2/ 5^2 = 169/25 sec = 13/5

OpenStudy (anonymous):

how can you change sec into sec^2x?

OpenStudy (anonymous):

by squaring :)

OpenStudy (anonymous):

13/5 is the right ans

OpenStudy (anonymous):

\[\sec x=\sqrt{1+\tan^2x}\]